OpenStudy (anonymous):

SOLVING THE PROPORTIONS??? Solve the proportions. 4. 5 / x + 2 = 3 / x + 1 5. (n-5)/(n+8) = 2 / 7 6. x / y - a / b ; solve for x.

OpenStudy (anonymous):

\[4. \frac{ 5 }{ x+2 } = \frac{ 3 }{ x + 1 }\] \[5. \frac{ (n-5) }{ (n+8) } = \frac{ 2 }{ 7 }\] 6. \[6. \frac{ x }{ y } = \frac{ a }{ b } ; solve for x.\]

OpenStudy (anonymous):

4. x = 1/2 5. n = 51/5 6. x = ay/b